Yesterday’s estimate of the daily variance rate of an asset was 1.44%2 and yesterday’s closing price of the asset was $50.00. Suppose that today’s closing price of the asset was $49.70. What is tomorrow’s estimate of the daily volatility of the asset using the EWMA model with the parameter \lambda of 0.94?
Answer #1

Using EWMA model,

V(t) = {Iambda*(V[ t-1]^2) +(1-lambda)*(current return^2)}^0.5

where lambda = 0.94; V[t-1] = 1.44%; current return = ln(today's price/yesterday's price) = ln(49.70/50) = -0.60181%

V(t) = {(0.94*1.44%^2) + (1-0.94)*(-0.60181%^2)}^0.5 = 1.40% (Tomorrow's estimate of daily volatility)